# BeeManc at the PLABA Track of TAC-2023: Investigating LLMs and Controllable Attributes for Improving Biomedical Text Readability

### v1

@@ 2.2 Automatic Evaluations
In this section, we list quantitative evaluation scores and some explanations for them. The results for T5 Small, T5 Base, T5 Large, FLAN-T5 XL with LORA, SciFive Base, SciFive Large, and BART models with CTs (BART-w-CTS) are displayed in Table 1. Interestingly, the fine-tuned T5 Small model obtains the highest scores in both BLEU and ROUGE metrics including ROUGE-1, ROUGE-2, and ROUGE-L. The fine-tuned BART Large with CTs (BART-L-w-CTs) produces the highest SARI score at 46.54; while the fine-tuned T5 Base model achieved the highest BERTScore (72.62) with a slightly lower SARI score (44.10). The fine-tuned SciFive Large achieved the highest SARI score (44.38) among T5-like models, though it is approximately 2 points lower than BART Large with CTs.
@@
The quantitative evaluation scores of GPT-like models are presented in Table 2 including GPT-3.5 and GPT-4 using prompts, and fine-tuned BioGPT with LoRA. GPT-3.5 reported relatively higher scores than GPT-4 on all lexical metrics except for SARI, and a much higher BERTScore than GPT-4 (58.35 vs 46.99). In comparison, BioGPT-Large with LoRA reported the lowest SARI score (18.44) and the highest BERTScore (62.9) among these three GPT-like models. Comparing the models across Table 1 and Table 2, the GPT-like models did not beat T5-Base on both SARI and BERTScore and did not beat BART-w-CTs on SARI. Therefore, we did not pick GPT-like models for our submission to the shared task, since each team can only submit up to three system outputs.
@@
To look into the details of model comparisons from different epochs on the extracted testing set, we present the learning curve of T5, SciFive, BART-base on WikiLarge (BART-b-Wiki), and continuously-learned BART-b-Wiki on the PLABA data (our training set only) in Figure 3. We also present the learning curve of T5 Base and BART Base using different metrics in Figure 4.
@@
Table 1: Automatic Evaluations of fine-tuned T5, SciFive, and BART Models with Control Token (BARTs-w-CTs) mechanisms on the Extracted Testing Set. FLAN-T5 XL used LoRA.
@@
Table 2: Automatic Evaluations of 1-shot GPTs and fine-tuned BioGPT-Large on the Extracted Testing Set
Figure 3: Learning Curves using BERTScores and SARI Scores of T5, SciFive and BART Models on the Extracted Testing Set
Figure 4: Learning Curves of T5-base and BART-base using BLEU, ROUGE, SARI, and BERTScores on the Extracted Testing Set
@@
Table 3: Sample of human evaluation questionnaire
@@
Table 4: Human Evaluation Scores of Fine-tuned BART-Large with CTs and T5-Base Models on the Sampled Test Set
@@
Table 5: Cohen Kappa among annotators over 3 categories ordinal - win, lose, and tie.
@@
Table 6: Krippendorff’s alpha among annotators (Anno.) over the 5-Likert scale from strongly agree to strongly disagree.
2.3 Human Evaluation
In our human evaluation, we randomly sampled 80 sentences from the test set split and evaluated the corresponding outputs of BART-large with CTs and T5-base in an anonymised setting. In the human evaluation form, we randomly assigned the order of the two systems’ outputs and put them beside the input sentence. Based on the comparison of input and output sentences, the annotators need to answer two questions: “To what extent do you agree the simplified sentence keeps the major information” and “To what extent do you agree the simplified sentence is well simplified”. The answer is limited to a 5-point Likert scale, from strongly disagree to strongly agree. The sample form can be found in Table 3.
There are 4 annotators in this human evaluation, one of the annotators is a native English speaker, and the others are fluent English users as a second language. Two annotators are final-year bachelor students, one annotator is a Masters candidate, and the last one is a Postdoctoral researcher. Each annotator evaluated 40 sentence pairs with 50% overlaps to make sure every sentence was evaluated twice by different annotators. The detailed human evaluation scores on the two selected models we evaluated are shown in Table 4 using the two designed criteria.
Based on the cross-annotation (overlaps), we calculated the inter-rater agreement level in Table 5 using Cohen’s Kappa on cross models and Table 6 using Krippendorff’s Alpha with model-wise comparison. Both tables include the agreement levels on two sub-categories, namely “meaning preservation” and “text simplicity”. Because there is no overlap in the annotation tasks between annotators (0, 3) and annotators (1, 2), we listed the agreement levels between the available pairs of annotations, i.e. (0, 1), (0, 2), (1, 3), and (2, 3). The Cohen’s Kappa agreement level presented in Table 5 shows the inter-rater agreement on the performance order of the two systems, whether one system is better than the other or tie. The Krippendorff’s Alpha represented in Table 6 shows the annotation reliability over all 5-Likert options.
Based on the results from Table 4, annotators show a different preference for the two systems. Despite the limited gap in the SARI score, BART with CTs shows a better capability to conduct the simplification tasks. Yet regarding meaning preservation, the fine-tuned T5-base performs better, as the BERTScore tells among the comparisons in Table 1.
From Table 5, Annotators 0 and 1 have the highest agreement on “meaning preservation” (score 0.583), while Annotators 1 and 3 have the highest agreement on “simplicity” (score 0.238) evaluation across the two models. This also indicates that it is not easy to evaluate the system performances against each other regarding these two criteria.
Model wise, Table 6 further shows that Annotators 0 and 1 agree more on the judgements of fine-tuned T5 Base model on both two criteria of “meaning preservation” (score 0.449) and “text simplicity” (score 0.386), in comparison to the BART model. On the fine-tuned BART Large model with CTs, these two annotators only agreed on the “meaning preservation” factor with a score of 0.441. This phenomenon also applies to Annotators 0 and 2 regarding the judgement of these two models. Interestingly, while Table 5 shows that Annotators 1 and 3 have a better agreement on “simplicity” judgement over “meaning preservation” using Cohen’s Kappa, Table 6 shows the opposite using Krippendorff’s Alpha, i.e. it tells the agreement on “meaning preservation” of these two annotators instead. This shows the difference between the two agreement and reliability measurement metrics.
3 BeeManc Submissions to PLABA2023
3.1 1st-round submissions
@@ 3 BeeManc Submissions to PLABA2023
For system submissions, we chose our domain and task fine-tuned BART-w-CTs as our first system, Biomedical-T5, and Lay-SciFive as our second and third ones.
@@
Figure 5: The official Human Evaluation on Simplicity, Accuracy, Fluency, Completeness, and Faithfulness at sentence-level, term-level, and averages. PLABA-base-1 was re-evaluated in round-2 submissions. Ondov et al. (2024)
Figure 6: Official SARI Scores from Teams computed against 4 references (submission round 1). BeeManc 1/2/3: BART-w-CTs, Biomedical-T5, Lay-SciFive Ondov et al. (2024)
3.2 2nd-round submissions
@@
4 Conclusions and Future Work
In this system description, we have reported our model development using LLMs and Control Mechanisms for the text simplification task on biomedical abstract using the PLABA data set and our submissions to PLABA-2023 track. From three of our submitted systems BART-w-CTs, Biomedical-T5, and Lay-SciFive, we have the following highlights in the official outcomes:
@@ 4 Conclusions and Future Work
In future work, we plan to carry out investigations on more recent models including BioBART Yuan et al. (2022), try different prompting methods such as the work from Cui et al. (2023), and design a more detailed human evaluation such as the work proposed by Gladkoff and Han (2022) with error severity levels might shed some light on this.
For this shared task challenge, we carried out data-constrained fine-tuning and model development only using the PLABA data set. However, in the broader sense, we would like to carry out data-augmented training such as using biomedical synthetic data generated by state-of-the-art Transformer-like models Belkadi et al. (2023).
Author Contributions
ZL carried out experiments on BART and Control Mechanisms, prepared Human Evaluation data, Measured inter-rater agreement. SB carried out experiments on T5 and SciFive. NM carried out experiments on GPTs and Bio-GPTs. LH supervised the project and helped drafting the system report. Human evaluation by ZL, SB, NM, and LH. Writing-wise, Abstract: LH; Section 1: ZL, SB, NM, LH; Section 2: ZL, SB, LH; Section 3: LH; Section 4: LH. Appendix: SB and LH. MS and GN: co-supervisors and approved the manuscript.
Acknowledgements
We thank XC for the valuable discussions, thank Brian Ondov and Hoa T. Dang for keeping in touch with us during the shared task. LH and GN are grateful for the support from the grant “Assembling the Data Jigsaw: Powering Robust Research on the Causes, Determinants and Outcomes of MSK Disease”. The project has been funded by the Nuffield Foundation, but the views expressed are those of the authors and not necessarily the Foundation. Visit www.nuffieldfoundation.org. LH and GN were also supported by the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).

### v2

@@ IV-B Automatic Evaluation Scores
In this section, we list quantitative evaluation scores and some explanations for them. The results for T5 Small, T5 Base, T5 Large, FLAN-T5 XL with LORA, SciFive Base, SciFive Large, and BART models with CTs (BART-w-CTS) are displayed in Table I. Interestingly, the fine-tuned T5 Small model obtains the highest scores in both BLEU and ROUGE metrics including ROUGE-1, ROUGE-2, and ROUGE-L. The fine-tuned BART Large with CTs produces the highest SARI score at 46.54; while the fine-tuned T5 Base model achieved the highest BERTScore (72.62) with a slightly lower SARI score (44.10). The fine-tuned SciFive Large achieved the highest SARI score (44.38) among T5-like models, though it is approximately 2 points lower than BART Large with CTs.
@@
The quantitative evaluation scores of GPT-like models are presented in Table II including GPT-3.5 and GPT-4 using prompts, and fine-tuned BioGPT with LoRA. GPT-3.5 reported relatively higher scores than GPT-4 on all lexical metrics except for SARI, and much higher score on BERTScore than GPT-4 (58.35 vs 46.99). In comparison, BioGPT-Large with LoRA reported the lowest SARI score (18.44) and the highest BERTScore (62.9) among these three GPT-like models. Comparing the models across Table I and Table II, the GPT-like models did not beat T5-Base on both SARI and BERTScore, and did not beat BART-w-CTs on SARI.
@@
To look into the details of model comparisons from different epochs on the extracted testing set, we present the learning curve of T5, SciFive, BART-base on WikiLarge, and BART-base on PLABA data in Figure 3. We also present the learning curve of T5 Base and BART Base using different metrics in Figure 4.
@@
TABLE I: Automatic Evaluations of Encoder-Decoder Models on Extracted Testing Set.
@@
TABLE II: Quantitative Evaluations of GPTs and BioGPT-Large on the Extracted Testing Set
Figure 3: Evaluation Scores of T5, SciFive and BART Models on the Extracted Testing Set
Figure 4: Evaluation Scores of T5-base and BART-base on the Extracted Testing Set
@@ IV-C Internal Human Evaluations
TABLE III: Sample of human evaluation questionnaire
In human evaluation, we randomly sampled 80 sentences from the test set split and evaluated the corresponding outputs of BART-large with CTs and T5-base with anonymisation setting. In the human evaluation form (Table III), we randomly assigned the order of the two systems’ outputs and put them beside the input sentence. Based on the comparison of input and output sentences, the annotators need to answer two questions: “To what extent do you agree the simplified sentence keeps the major information" and “To what extent do you agree the simplified sentence is well simplified". The answer is limited to a 5-point Likert scale, from strongly disagree to strongly agree. There are 4 annotators in this human evaluation, one of the annotators is a native English speaker, and the others are fluent English users as a second language. Two annotators are final-year bachelor students, one annotator is a Masters candidate, and the last one is a Postdoctoral researcher. Each annotator evaluated 40 sentence pairs with 50% overlaps to make sure every sentence was evaluated twice by different annotators. The detailed human evaluation scores on the two selected models we evaluated are shown in Table IV using the two designed criteria.
Based on the cross-annotation (overlaps), we calculated the inter-rater agreement level in Table V using Cohen’s Kappa on cross models and Table VI using Krippendorff’s Alpha with model-wise comparison. Both tables include the agreement levels on two sub-categories, namely “meaning preservation" and “text simplicity". Because there is no overlap in the annotation tasks between annotators (0, 3) and annotators (1, 2), we listed the agreement levels between the available pairs of annotations, i.e. (0, 1), (0, 2), (1, 3), and (2, 3). The Cohen’s Kappa agreement level presented in Table V shows the inter-rater agreement on the performance order of the two systems, whether one system is better than the other or tie. The Krippendorff’s Alpha represented in Table VI shows the annotation reliability over all 5-Likert options.
Based on the results from Table IV, annotators show a different preference for the two systems. Despite the limited gap in the SARI score, BART with CTs shows a better capability to fulfil the simplification tasks. Yet regarding meaning preservation, the fine-tuned T5-base performs better, as the BERTScore tells among the comparisons in Table I.
From Table V, Annotators 0 and 1 have the highest agreement on “meaning preservation” (score 0.583), while Annotators 1 and 3 have the highest agreement on “simplicity” (score 0.238) evaluation across the two models. This also indicates that it is not easy to evaluate the system performances against each other regarding these two criteria.
Model wise, Table VI further shows that Annotators 0 and 1 agree more on the judgements of fine-tuned T5 Base model on both two criteria of “meaning preservation” (score 0.449) and “text simplicity” (score 0.386), in comparison to the BART model. On the fine-tuned BART Large model with CTs, these two annotators only agreed on the “meaning preservation” factor with a score of 0.441. This phenomenon also applies to Annotators 0 and 2 regarding the judgement of these two models. Interestingly, while Table V shows that Annotators 1 and 3 have a better agreement on “simplicity” judgement over “meaning preservation” using Cohen’s Kappa, Table VI shows the opposite using Krippendorff’s Alpha, i.e. it tells the agreement on “meaning preservation” of these two annotators instead. This shows the difference between the two agreement and reliability measurement metrics.
@@
TABLE IV: Human Evaluation Scores of Fine-tuned BART-Large with CTs and T5-Base Models on the Sampled Test Set
@@
TABLE V: Cohen Kappa among annotators over 3 categories ordinal - win, lose, and tie.
@@
TABLE VI: Krippendorff’s alpha among annotators (Anno.) over the 5-Likert scale from strongly agree to strongly disagree.
IV-D System Output Categorisation
We observe some interesting aspects of human evaluation findings by comparing the outputs of two models. 1) how to deal with the judgement on two models when one almost copied the full text from the source while the other did simplification but with introduced errors? 2) Abbreviation caused interpretation inaccuracy. This can be a common issue in the PLABA task. For instance, the source sentence “A total of 157 consecutive patients underwent TKA (n = 18) or UKA (n = 139).” is simplified by the T5-base model into “A total of 157 consecutive patients underwent knee replacement or knee replacement.” and by BART-w-CTs into “A total of 157 patients had either knee replacement or knee replacement surgery.” Both these two models produced repeated phrases “knee replacement or knee replacement” due to a lack of meaningful understanding of “TKA: total knee arthroplasty” and “UKA: Unicompartmental knee arthroplasty”. A reasonable simplification here can be “157 patients had knee surgery”.
We list more categories below by comparing outputs and refer to Table LABEL:tab:human_eval_catergorise for examples from the two evaluated models. These categories apply to some segment-level comparisons, but not all the outputs, e.g. in Figure 5 and 6. From the example outputs, we can see — “Large Language Models for Biomedical Text Simplification: Promising But Not There Yet”, as stated in our paper title.
  • Both models simplified the abstract into the exact same output
  • Both models produced hallucinations and similar outputs
  • Both models cut half the sentence/meaning, but at different parts
  • Both models cut complex sentences into multiple sentences, but BART adapted lay-language
  • BART adapted using lay language but cut out some meaning
  • BART generated simplification vs T5 generated nonsense
  • T5 did little simplification but maintained good meaning.
  • T5 cut meaning; BART did not but maintained the same complexity as the abstract
  • BART shifted the meaning
Figure 5: Segment-level examples of categorised outputs from two models (Part-1).
Figure 6: Segment-level examples of categorised outputs from two models (Part-2).
V Discussions
Several aspects can be further improved or investigated:
  • On automatic metrics. Based on the results depicted in Figure 4, we acknowledge that SARI stands out as a more reliable metric than BLEU and ROUGE-1/2 for assessing the quality of generated simplifications. During the early training epochs, the model outputs closely resemble the input texts, which can lead SARI to assign lower scores compared to BLEU and ROUGE-1/2. This occurs because these metrics might be satisfied by the mirroring of n-grams between inputs and generated outputs during simplification. However, it is essential to acknowledge that, in the context of simplification, the generated output typically remains relatively close to the input. As a result, BLEU and ROUGE may exhibit consistent scores throughout epochs and may not effectively evaluate the quality of the generated texts. In contrast, BERTScore offers a different perspective by focusing on meaning preservation after simplification instead of simplification quality. If the generated outputs are copies of the input texts, BERTScore may still yield high scores although the model indeed performs poorly. Therefore, we use the combination of both metrics — SARI for evaluating generation quality and BERTScore for assessing meaning preservation — in order to select the best-performing models. It would be very useful to develop a new automatic metric that can effectively reflect both text simplicity and meaning preservation.
  • On internal human evaluations. Due to the various backgrounds of annotators and the lack of proper training material and methods, it is difficult to build a standardised scheme to help annotators choose from the 5-Likert options. This means that the definition of “Strongly Agree/Disagree” may vary among annotators. In addition, there is no method to normalise the scores to a unified level in order to avoid the effects caused by differences in subjective judgements. Thus, we evaluate the two systems simultaneously and allow the annotator to decide on the performance order and performance gap. To reflect on the agreement of these two aspects, we calculated the inter-rater agreement in Cohen Kappa and Krippendorff’s alpha. As shown in Table V, only annotators 0 and 1 show a decent agreement on meaning preservation, while the others show only limited agreement or even slight disagreement. Although human evaluation has long been the gold standard for evaluation tasks, there are few acknowledged works on a standard procedure to make it more explainable and comparable. In the future, a more standardised and unified process may be required.
  • On broader test sets. In this work, we carried out an investigation using the PLABA data set. To be more fairly comparing selected models, we will include other related testing data on biomedical text simplification tasks. How to improve the number of references for the PLABA data set can be also explored.
VI BeeManc Submissions to PLABA-2023
VI-A Round-1: BART-w-CT, Biomedical-T5, Lay-SciFive
@@
For system submissions, in Round-1, we chose our domain and task fine-tuned BART-w-CTs as our first system, Biomedical-T5, and Lay-SciFive as our second and third ones.
@@
Figure 7: The official Human Evaluation on Simplicity, Accuracy, Fluency, Completeness, and Faithfulness at sentence-level, term-level, and averages. PLABA-base-1 was re-evaluated in round-2 submissions. [57]
Figure 8: Official SARI Scores from Teams computed against 4 references (submission round 1). BeeManc 1/2/3: BART-w-CTs, Biomedical-T5, Lay-SciFive [57]
VI-B Round-2: OpenAI’s GPTs
Figure 9: GPT Prompt Examples
@@ VI-B Round-2: OpenAI’s GPTs
VII Conclusions and Future Work
We have carried out an investigation into using LLMs and Control Mechanisms for the text simplification task on biomedical abstracts using the PLABA data set. Both automatic evaluations using a broad range of metrics and human evaluations were conducted to assess the system outputs. As our automatic evaluation results show, both T5 and BART with Control Tokens demonstrated high accuracy in generating simplified versions of biomedical abstracts. However, when we delve into human evaluations, it becomes clear that each model possesses its unique strengths and trade-offs. T5 demonstrated strong performances at preserving the original abstracts’ meaning, but sometimes at the cost of lacking simplification. By maintaining the core content and context of the input, it has proven to be over-conservative in some cases, resulting in outputs that very closely resemble the inputs therefore maintaining the abstract’s complexity. On the other hand, BART-w-CTs demonstrated strong simplification performances to produce better-simplified versions. However, it has shown a potential drawback in reducing the preservation of the original meaning.
We also reported our BeeManc team submission outcomes to the official PLABA 2023 challenges. From three of our submitted systems BART-w-CTs, Biomedical-T5, and Lay-SciFive, we have the following highlights in the official outcomes:
@@ VII Conclusions and Future Work
For this shared task challenge, we carried out data-constrained fine-tuning and model development only using the PLABA data set. However, in the broader sense, we would like to carry out data-augmented training such as using biomedical synthetic data generated by state-of-the-art Transformer-like models [58].
In future work, we plan to carry out investigations on more recent models including BioBART [25], try different prompting methods such as the work from [59], and design a more detailed human evaluation such as the work proposed by [60] with error severity levels might shed some light on this.